OpenStudy (anonymous):

does anyone know 5/6+4/9=?

OpenStudy (anonymous):

First you have to make the denominators the same. To do that, you multiply the numbers by different forms of 1, such as 5/5 or 3/3, which still leaves them equivalent.

OpenStudy (anonymous):

So do (5/6*3/3)+(4/9*2/2)

OpenStudy (anonymous):

I got 1.27777777778 and I hope this helped:)

OpenStudy (anonymous):

which is 15/18+8/18=23/18 If you teacher is picky, then you can change it to a mixed number which is \[1\frac{ 7 }{ 18 }\]

OpenStudy (anonymous):

are u look for the LCD?

OpenStudy (anonymous):

That is how you do the problem lol

OpenStudy (anonymous):

ok, thank u guys for helping :)

OpenStudy (♪chibiterasu):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Rubyjadeluv np and yea the LCD is 40 \(\color{#0cbb34}{\text{End of Quote}}\) The LCD is 18 not 40. 40 is not divisble by 6 nor 9.

OpenStudy (anonymous):

18?

OpenStudy (♪chibiterasu):

I don't even know where the 40 came from.

OpenStudy (triciaal):

The "proper" way to find the lcd is to find the prime factors then only use the common factors once 6 :2 *3 9: 3*3 3 is common only use once. Lcd = 2*3*3 = 18